The temperature at the center of sun is 1.55×107k.convert this kelvin temperature to the celsius and fahrenheit scales

Grade:12

1 Answers

Nirmal Singh.
10 years ago
we know that Kelvin = 273 + celsius
Celsius = Kelvin - 273 = (1.55x10^7 - 273) C
Also we know (F -32)/ 9 = C /5
so F = 9/5C +32 = 9/5(1.55x10^7 - 273) +32 F
Thanks & Regards,
Nirmal Singh
Askiitians Faculty